1. Using Snell’s law, determine the constant when the angle of incidence and the angle of refraction are: a) 50˚ and 30˚ b) 30˚ and 18˚ c) 60˚ and 38˚


2. a) What is the angle of refraction in a medium if the angle of incidence in air is 48˚ and the index of refraction of the medium is 1.58? b) What is the angle of incidence in a medium in the case where the angle of refraction in air is 40˚ and the index of refraction of the medium is 1.58?


3. A ray of light passes from air into water at an angle of incidence of 50˚. What is the angle of refraction? (HINT: You will need to use the sin1 function on your calculator).


1
Expert's answer
2022-05-30T15:17:17-0400

"n=\frac{\sin\alpha}{\sin\beta},"

a)

"n_1=1.53,"

b)

"n_2=1.62,"

c)

"n_3=1.41."